OpenStudy (anonymous):

Simplify the following expression and write the answer in simplest form. (5/4)^2 + 7/16 A. 1 1/16 B.32/16 C. none of these D. 2 I know its B but can it be simplified?

OpenStudy (anonymous):

Yes it can be simplified, 32/16= 2. Which would mean your answer would be D.2

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{5}{4}\right)^2+\cfrac{7}{16}\implies \cfrac{5^2}{4^2}+\cfrac{7}{16}\implies \cfrac{25}{16}+\cfrac{7}{16}=\cfrac{}{LCD?}\)

OpenStudy (texaschic101):

(5/4)(5/4) + 7/16 25/16 + 7/16 = you can do this

OpenStudy (anonymous):

@jdoe0001 32/16?

OpenStudy (jdoe0001):

welll.. actually you got correct... is= 32/16 yes notice that 16* 2 = ?

OpenStudy (texaschic101):

yep...32/16

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{5}{4}\right)^2+\cfrac{7}{16}\implies \cfrac{5^2}{4^2}+\cfrac{7}{16}\implies \cfrac{25}{16}+\cfrac{7}{16}=\cfrac{\cancel{ 32 }}{\cancel{ 16 }}\to \cfrac{2}{1}\to 2\)

OpenStudy (texaschic101):

32/16 = 8/4 = 2...your right, it can be simplified to 2

OpenStudy (jdoe0001):

so one can say that B or D are both correct

OpenStudy (anonymous):

thanks for your help guys! :D

OpenStudy (texaschic101):

I agree...both are correct

OpenStudy (jdoe0001):

but the "simplest form" would be 2

OpenStudy (texaschic101):

true

OpenStudy (jdoe0001):

yw